OpenStudy (anonymous):

Check my answer please! I will give a medal! Simplify the expression. y + 1.2y + 1.2z My answer: y = 1y 1y + 1.2y = 2.2y Final answer: 2.2y + 1.2z

OpenStudy (anonymous):

I would agree with your answer. You seem to know your stuff XD I'm sure you can understand why it is right. If you don't let me know and I can exaplin further.

OpenStudy (whpalmer4):

Seems reasonable to me.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

That is correct. y + 1.2y + 1.2z = (1 + 1.2)y + 1.2z =2.2y+1.2z